as shown in fig. 5-14, the uniform, 1600-n beam is hinged at one end and held try a tic cope at the other

1
Expert's answer
2022-01-17T09:59:32-0500

"T=\frac N{\cos\alpha},"

"T=\frac{1600}{0.77}=2100~N."